# Designing Phase Sensitive Probes of Monopole Superconducting Order

### v1

@@ Appendix C Monopole harmonic functions
Monopole harmonics are eigenfunctions of angular momentum operator in the presence of a magnetic monopole [11, 37]. Due to the magnetic monopole, the vector potential $\mathbf{A}$ must have singularities, corresponding to a Dirac string. To avoid these singularities, one partitions the unit sphere into two regions covering the north and south poles. For a magnetic monopole of strength $g$, the vector potential is given by
@@
in which the superscript $\lambda=\pm 1$ corresponds to a gauge in which the vector potential is well-defined, with the Dirac string located at $\theta=(\lambda-1)\pi/2$. For a particle of electric charge $e$ moving in the presence of a magnetic monopole, let $q={ge}/{(\hbar c)}={n}/{2}$ for $n\in\mathbb{Z}$ denote the monopole charge.
@@
where $q$ is the corresponding monopole charge, which sets the lower bound for the partial wave channels, $l\geq q$. Above, $\tilde{\mathbf{L}}={\mathbf{r}}\times({\mathbf{p}}+|e|\mathbf{A}^{(\lambda)}/\hbar c)-{\lambda}\hbar q\hat{\mathbf{r}}$ and $\tilde{L}_{z}=-i\hbar\partial_{\varphi}-{\lambda}\hbar q=L_{z}^{(\lambda)}-{\lambda}\hbar q$. Due to the nontrivial monopole charge, functions defined in the $\lambda=+1$ and $\lambda=-1$ patch are related by a $\mathrm{U}(1)$ transformation,
@@
in which the superscript corresponds to a gauge choice consistent with Eq. (70). In the following, the monopole harmonics are normalized to the convention $\int_{0}^{-1}\mathrm{d}(\cos\theta)\int_{0}^{2\pi}\mathrm{d}\varphi|\mathcal{Y}_{q,l,l_{z}}(\theta,\varphi)|^{2}=1$.
The general form of the monopole harmonic $\mathcal{Y}_{q;l,l_{z}}$ is given by [11]
@@
in which $\alpha=-q-l_{z}$, $\beta=q-l_{z}$, $n=j+l_{z}$, $x=\cos\theta$, $M_{q,l,l_{z}}=2^{l_{z}}\sqrt{{(2l+1)}/{4\pi}}\sqrt{{(l-l_{z})!(l+l_{z})!}/{(l-q)!(l+q)!}},$ and $P_{n}^{\alpha,\beta}$ are Jacobi polynomials,
$P_{n}^{\alpha,\beta}(x)=\frac{(-1)^{n}}{2^{n}n!}(1-x)^{-\alpha}(1+x)^{-\beta}\frac{\mathrm{d}^{n}}{\mathrm{d}x^{n}}\left[(1-x)^{\alpha+n}(1+x)^{\beta+n}\right].$ (74)
@@
Monopole harmonics of opposite charge are related by $\mathcal{Y}_{q;l,l_{z}}(\theta,\varphi)=(-1)^{q+l_{z}}\mathcal{Y}^{*}_{-q;l,-l_{z}}(\theta,\varphi).$ The product of two monopole harmonics can be expressed as [37]
@@
in which $q_{3}=q_{1}+q_{2}$ and $l_{3,z}=l_{1,z}+l_{2,z}$, and the round brackets are the Wigner $3j$ symbols. In other words, the product of two monopole harmonics can be expressed as a superposition of monopole harmonics, all which belong in the same topological sector $q_{3}$ and have global angular momentum $l_{3,z}$, but may reside in different partial wave channels.
@@
For $q=-1$, the monopole harmonics in the $l=1,2$ partial wave channel read:
q = -$\begin{array}[]{c|c|c}\textbf{q = -1}&l=1&l=2\\ \hline\cr l_{z}=-2&&\sqrt{\frac{5}{4\pi}}\sin^{2}\frac{\theta}{2}\sin\theta e^{-i(\lambda+2)\varphi}\\ \hline\cr l_{z}=-1&\sqrt{\frac{3}{4\pi}}\sin^{2}\frac{\theta}{2}e^{-i(\lambda+1)\varphi}&\sqrt{\frac{5}{4\pi}}\sin^{2}\frac{\theta}{2}(1+2\cos\theta)e^{-i(\lambda+1)\varphi}\\ \hline\cr l_{z}=0&\sqrt{\frac{3}{8\pi}}\sin\theta e^{-i\lambda\varphi}&\sqrt{\frac{15}{8\pi}}\cos\theta\sin\theta e^{-i\lambda\varphi}\\ \hline\cr l_{z}=1&\sqrt{\frac{3}{4\pi}}\cos^{2}\frac{\theta}{2}e^{i(-\lambda+1)\varphi}&\sqrt{\frac{5}{4\pi}}\cos^{2}\frac{\theta}{2}(-1+2\cos\theta)e^{i(-\lambda+1)\varphi}\\ \hline\cr l_{z}=2&&\sqrt{\frac{5}{4\pi}}\cos^{2}\frac{\theta}{2}\sin\theta e^{i(-\lambda+2)\varphi}\end{array}$
@@ Appendix D Tight-binding model of Josephson junction
We describe the tight-binding model used to model the SIS Josephson junction, as used in Sec. IV.1.2, V, and VI. For the following analysis, we consider a junction along the $x$ direction without loss of generality. We take periodic boundary conditions in the $y$ and $z$ directions, with transverse momenta $k_{y}$ and $k_{z}$ being conserved across the junction barrier. For given conserved momenta $k_{y}$ and $k_{z}$, we treat the SIS junction as a pseudo-1D junction in the $x$ direction with $2N_{x}$ sites, in which the barrier is located between $n_{x}=N_{x}$ and $n_{x}=N_{x}+1$. We consider the Hamiltonian of a Josephson junction, $H_{\mathrm{JJ}}(k_{y},k_{z})=H_{\mathrm{BdG},L}(k_{y},k_{z})+H_{\mathrm{BdG},\mathrm{link}}(k_{y},k_{z})+H_{\mathrm{BdG},L}(k_{y},k_{z})$ in Eq. (1). The bulk superconductors at the left (right) side are given by
@@
in which sum over $n_{x}$ runs from $1$ to $N_{x}$ for the left superconductor ($N_{x}+1$ to $2N_{x}$ for the right superconductor). Above, $\mathcal{H}_{\mathrm{kin},L(R)}(k_{y},k_{z})$ is the kinetic Hamiltonian, $\Delta_{L(R)}(k_{y},k_{z})$ is the superconducting pairing order parameter, and $\phi_{L(R)}$ is the $\mathrm{U}(1)$ superconducting phase, taken to be uniform in the respective superconductor. $c_{k_{y},k_{z};\sigma,n_{x}}$ is the annihilation operator for an electron at site $n_{x}$ with spin $\sigma$ and conserved momentum $k_{y}\hat{y}+k_{z}\hat{z}$.
The tight-binding analogue of the tunneling Hamiltonian for the SIS junction in Eq. (3) is given by
@@
in which $\hat{t}=t_{0}(k_{y},k_{z})\sigma_{0}+\mathbf{t}(k_{y},k_{z})\cdot\bm{\sigma}$ describes spin-independent and spin-dependent hopping across the junction barrier. Here, we consider only one site at the barrier for simplicity. Including more sites will increase the amount of in-gap Andreev states but not affect the overall qualitative features of the Josephson current, e.g. its periodicity.
@@ Appendix E Tight-binding models of monopole superconductors
In this section, we outline the tight-binding model of the monopole superconductor described in Eq. (13) used for numerical calculations of monopole superconductors. The general Hamiltonian of a monopole superconductor reads
$H_{\text{MSC}}=$ $\sum_{\mathbf{n},\sigma;\mathbf{n}^{\prime},\sigma^{\prime}}c^{\dagger}_{\mathbf{n},\sigma}[\mathcal{H}_{\mathrm{kin}}]_{\mathbf{n},\sigma;\mathbf{n}^{\prime},\sigma^{\prime}}c_{\mathbf{n}^{\prime},\sigma^{\prime}}+\sum_{\mathbf{n},\sigma;\mathbf{n}^{\prime},\sigma^{\prime}}\left(c^{\dagger}_{\mathbf{n},\sigma}[\Delta_{\mathrm{inter}}]_{\mathbf{n},\sigma;\mathbf{n}^{\prime},\sigma^{\prime}}c^{\dagger}_{\mathbf{n}^{\prime},\sigma^{\prime}}+\mathrm{h.c.}\right)$ (79)
in which $\mathcal{H}_{\mathrm{kin}}$ and $\Delta_{\mathrm{inter}}$ are the kinetic and inter-Fermi-surface pairing components of the BdG Hamiltonian respectively in the combined real-space and spin representation. Above, $c_{\mathbf{n},\sigma}$ annihilates a single-particle state at lattice site $\mathbf{n}$ and with spin $\sigma=\uparrow,\downarrow$. $[\mathcal{H}_{\mathrm{kin}}]_{\mathbf{n},\sigma;\mathbf{n}^{\prime},\sigma^{\prime}}$ and $[\Delta_{\mathrm{inter}}]_{\mathbf{n},\sigma;\mathbf{n}^{\prime},\sigma^{\prime}}$ can be organized into nested blocks, with each block being a $2\times 2$ matrix with indices $\sigma$ and $\sigma^{\prime}$ denoting the spin degree of freedom. The blocks are denoted by lattice indices $(\mathbf{n},\mathbf{n}^{\prime})$, corresponding to the matrix elements between states localized at the sites $\mathbf{n}$ and $\mathbf{n}^{\prime}$.
E.1 $q_{p}=-1$ monopole superconductor
We provide the tight-binding model of the monopole superconductor with pair monopole charge $q_{p}=-1$, for which the continuum model is given in Eq. (18). The Fourier transform of the tight-binding model of the magnetic doped Weyl semimetal band Hamiltonian on a cubic lattice is given by
$\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}(\mathbf{k})=t_{z}(\cos K_{0,z}-\cos k_{z})\sigma_{z}+m(\cos k_{x}+\cos k_{y}-2)\sigma_{z}+t_{x}\sin k_{x}\sigma_{x}+t_{y}\sin k_{y}\sigma_{y}-\mu\sigma_{0},$ (80)
in which $\mu$ is the chemical potential and $t_{i=x,y,z}$ are the hopping amplitudes along the $i^{\text{th}}$ direction (we take $t_{x}=t_{y}=t_{z}\sin K_{0,z}=v_{F}$ to compare to Eq. (19)). At appropriate doping, the system gives rise to two disjoint Fermi surfaces surrounding Weyl nodes at ±$\pm\mathbf{K}_{0}=(0,0,\pm K_{0,z})^{\mathrm{T}}$. The parameter $m$ is chosen satisfying $m/t_{z}\leq-2$ so that the sign of the $\sigma_{z}$-term in momentum space only changes between the two Weyl nodes at ±$\pm\mathbf{K}_{0}$.
@@ E.1 $q_{p}=-1$ monopole superconductor
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n};\mathbf{n}}=(t_{z}\cos K_{0,z}-2m)\sigma_{z}-\mu\sigma_{0},$ (81a)
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{x}}=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n}+\bm{\delta}_{x};\mathbf{n}}^{\dagger}=\frac{1}{2}(m\sigma_{z}-it_{x}\sigma_{x}),$ (81b)
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{y}}=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n}+\bm{\delta}_{y};\mathbf{n}}^{\dagger}=\frac{1}{2}(m\sigma_{z}-it_{y}\sigma_{y}),$ (81c)
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{z}}=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n}+\bm{\delta}_{z};\mathbf{n}}^{\dagger}=-\frac{1}{2}t_{z}\sigma_{z}.$ (81d)
Above, $\bm{\delta}_{x}=(1,0,0)^{\mathrm{T}}$, $\bm{\delta}_{y}=(0,1,0)^{\mathrm{T}}$, and $\bm{\delta}_{z}=(0,0,1)^{\mathrm{T}}$ correspond to translations by one lattice vector along the crystalline axes, with the lattice constant taken to be unity. The inter-Fermi surface $s$-wave pairing is given by
@@
in which $\Delta_{0}$ is the pairing amplitude.
@@
Under open boundary conditions, the system exhibits helical surface states consistent with the local angular momentum, $l_{z,\mathrm{loc}}$. We consider the partial Fourier-transform of the above tight-binding model, such that $\tilde{k}_{z}$ is a conserved quantity. For given conserved $\tilde{k}_{z}$, the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor can be regarded as an effective two-dimensional chiral superconductor characterized by local angular momentum $l_{z,\mathrm{loc}}=\pm 1$. Away from $\mathrm{FS}_{1}$, the chiral surface modes result from the Weyl Fermi arcs; however, between the gap nodes at the poles of $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$, a single surface mode appears within the gap. The single surface mode appearing between the gap nodes is a result of the effective local chiral $p$-wave pairing, as described in Sec. V. As $\tilde{k}_{z}$ varies from the south to north pole, corresponding to Fig. 8 (b-d), though the topology of the pairing order, which is given by pair monopole charge $q_{p}=-1$, does not change, the sign of the in-gap states’ group velocity changes. The change in the group velocity is due to the shift in the local angular momentum, in agreement to the effective pairing order in Eq. (24).
@@
Figure 8: Surface modes of the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor solved under semi-infinite boundary conditions in the $x$ direction, computed with the tight-binding model. (a) Fermi surface $\mathrm{FS}_{1}$ surrounding the Weyl node at $\mathbf{K}_{0}$ (black, solid). (b-d) Bulk (gray) and surface spectra (colored) shown at momenta $\tilde{k}_{z}=-0.20$, $0.00$, and $0.17$ relative to the Weyl node, corresponding to the dashed lines in (a). Localized in-gap surface states are colored on a continuum with respect to their charge, with particle (hole) states corresponding to red (blue). As $k_{z}$ is varied between the gap nodes, the group velocity of the Majorana surface mode changes sign. Results are computed with $t_{x}=t_{y}=t_{z}\sin K_{0}=-0.5$, $K_{0}=1.0$, $\mu=-0.1$, $m=1.0$, and $\Delta_{0}=0.02$.
@@
We additionally examine the effective pairing projected to the Fermi surface of the tight-binding model. The effective pairing between two Fermi surfaces $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$, as introduced in Sec. III.1, is defined as $\Delta_{\mathrm{MSC}}^{(q_{p})}(\tilde{\mathbf{k}})=\langle\chi_{1,+}(\tilde{\mathbf{k}})|\Delta_{\mathrm{inter}}(\mathbf{k})|\chi_{2,+}^{*}(-\tilde{\mathbf{k}})\rangle$ in Eq. (16). Here, $|\chi_{1(2),+}(\tilde{\mathbf{k}})\rangle$ is the helical eigenstate of $\mathrm{FS}_{1(2)}$. In Fig. 9, we present the effective pairing computed with and projected onto $\mathrm{FS}_{1}$ of the tight-binding model of the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor outlined above. The magnitude of the effective pairing is shown in Fig. 9 (a-b), in which there are two gap nodes pinned to the north and south pole of $\mathrm{FS}_{1}$. The phase winding, chosen in a gauge in which the pairing analytic at the north and south pole, is shown in Fig. 9 (c) and (d) respectively. Near the south pole, analogous to the surface modes in Fig. 8 (b), the superconductor behaves locally as a $p_{x}+ip_{y}$ superconductor. In contrast, near the north pole, the effective pairing order behaves locally as a $p_{x}-ip_{y}$ superconductor. The shift in local angular momentum $l_{z,\mathrm{loc}}=\pm 1$ from the global angular momentum, $l_{z,\mathrm{glob}}=0$, is indicative of the nontrivial pair monopole charge, $q_{p}=-1$. For the tight-binding model, the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconducting pairing order can be expressed as a superposition of monopole harmonics, all which have pair monopole charge $q_{p}=-1$ and global angular momentum $l_{z}=2$, but reside in different partial wave channels ($l\geq|q_{p}|$) due to the lattice symmetry.

### v2

@@ Appendix C Monopole harmonic functions
Monopole harmonics are eigenfunctions of angular momentum operator in the presence of a magnetic monopole Wu and Yang (1976, 1977). Due to the magnetic monopole, the vector potential $\mathbf{A}$ must have singularities, corresponding to a Dirac string. To avoid these singularities, one partitions the unit sphere into two regions covering the north and south poles. For a magnetic monopole of strength $g$, the vector potential is given by
@@
where the superscript $\lambda=\pm 1$ corresponds to a gauge where the vector potential is well-defined, with the Dirac string located at $\theta=(\lambda-1)\pi/2$. For a particle of electric charge $e$ moving in the presence of a magnetic monopole, let $q={ge}/{(\hbar c)}={n}/{2}$ for $n\in\mathbb{Z}$ denote the monopole charge.
@@
where $q$ is the corresponding monopole charge, which sets the lower bound for the partial wave channels, $l\geq q$. Here, $\tilde{\mathbf{L}}={\mathbf{r}}\times({\mathbf{p}}+|e|\mathbf{A}^{(\lambda)}/\hbar c)-{\lambda}\hbar q\hat{\mathbf{r}}$ and $\tilde{L}_{z}=-i\hbar\partial_{\varphi}-{\lambda}\hbar q=L_{z}^{(\lambda)}-{\lambda}\hbar q$. Due to the nontrivial monopole charge, functions defined in the $\lambda=+1$ and $\lambda=-1$ patch are related by a $\mathrm{U}(1)$ transformation,
@@
where the superscript corresponds to a gauge choice consistent with Eq. (71). In the following, the monopole harmonics are normalized to the convention $\int_{0}^{-1}\mathrm{d}(\cos\theta)\int_{0}^{2\pi}\mathrm{d}\varphi|\mathcal{Y}_{q,l,l_{z}}(\theta,\varphi)|^{2}=1$.
The general form of the monopole harmonic $\mathcal{Y}_{q;l,l_{z}}$ is given by Wu and Yang (1976)
@@
where $\alpha=-q-l_{z}$, $\beta=q-l_{z}$, $n=j+l_{z}$, $x=\cos\theta$, $M_{q,l,l_{z}}=2^{l_{z}}\sqrt{{(2l+1)}/{4\pi}}\sqrt{{(l-l_{z})!(l+l_{z})!}/{(l-q)!(l+q)!}},$ and $P_{n}^{\alpha,\beta}$ are Jacobi polynomials,
$P_{n}^{\alpha,\beta}(x)=$ $\frac{(-1)^{n}}{2^{n}n!}(1-x)^{-\alpha}(1+x)^{-\beta}$ (75)
×$\times\frac{\mathrm{d}^{n}}{\mathrm{d}x^{n}}\left[(1-x)^{\alpha+n}(1+x)^{\beta+n}\right].$
@@
Monopole harmonics of opposite charges are related by $\mathcal{Y}_{q;l,l_{z}}(\theta,\varphi)=(-1)^{q+l_{z}}\mathcal{Y}^{*}_{-q;l,-l_{z}}(\theta,\varphi).$ The product of two monopole harmonics follows Wu and Yang (1977)
@@
where $q_{3}=q_{1}+q_{2}$ and $l_{3,z}=l_{1,z}+l_{2,z}$, and the round brackets are the Wigner $3j$ symbols. In other words, the product of two monopole harmonics can be expressed as a superposition of monopole harmonics, all which belong in the same topological sector $q_{3}$ and have global angular momentum $l_{3,z}$, but may reside in different partial wave channels.
@@
For $q=-1$, the monopole harmonics in the $l=1,2$ partial wave channel take the form:
q = -$\begin{array}[]{c|c}\textbf{q = -1}&l=1\\ \hline\cr l_{z}=-1&\sqrt{\frac{3}{4\pi}}\sin^{2}\frac{\theta}{2}e^{-i(\lambda+1)\varphi}\\ \hline\cr l_{z}=0&\sqrt{\frac{3}{8\pi}}\sin\theta e^{-i\lambda\varphi}\\ \hline\cr l_{z}=1&\sqrt{\frac{3}{4\pi}}\cos^{2}\frac{\theta}{2}e^{i(-\lambda+1)\varphi}\end{array}$
and
q = -$\begin{array}[]{c|c}\textbf{q = -1}&l=2\\ \hline\cr l_{z}=-2&\sqrt{\frac{5}{4\pi}}\sin^{2}\frac{\theta}{2}\sin\theta e^{-i(\lambda+2)\varphi}\\ \hline\cr l_{z}=-1&\sqrt{\frac{5}{4\pi}}\sin^{2}\frac{\theta}{2}(1+2\cos\theta)e^{-i(\lambda+1)\varphi}\\ \hline\cr l_{z}=0&\sqrt{\frac{15}{8\pi}}\cos\theta\sin\theta e^{-i\lambda\varphi}\\ \hline\cr l_{z}=1&\sqrt{\frac{5}{4\pi}}\cos^{2}\frac{\theta}{2}(-1+2\cos\theta)e^{i(-\lambda+1)\varphi}\\ \hline\cr l_{z}=2&\sqrt{\frac{5}{4\pi}}\cos^{2}\frac{\theta}{2}\sin\theta e^{i(-\lambda+2)\varphi}\end{array}$
@@ Appendix D Tight-binding model of Josephson junction
We describe the tight-binding model used to model a Josephson junction, as used in Sec. IV.1.2, V, and VII. For the following analysis, we consider a junction along the $x$ direction without loss of generality. We take periodic boundary conditions in the $y$ and $z$ directions, with transverse momenta $k_{y}$ and $k_{z}$ being conserved across the junction barrier. For given conserved momenta $k_{y}$ and $k_{z}$, we treat the junction as a pseudo-$1$D junction in the $x$ direction with $2N_{x}$ sites, where the barrier is located between $n_{x}=N_{x}$ and $n_{x}=N_{x}+1$. We consider the Hamiltonian of a Josephson junction, $H_{\mathrm{JJ}}(k_{y},k_{z})=H_{\mathrm{BdG},L}(k_{y},k_{z})+H_{\mathrm{link}}(k_{y},k_{z})+H_{\mathrm{BdG},R}(k_{y},k_{z})$ in Eq. (1). The bulk superconductors at the left (right) side are given by
@@
where sum over $n_{x}$ runs from $1$ to $N_{x}$ for the left superconductor ($N_{x}+1$ to $2N_{x}$ for the right superconductor). Here, $\mathcal{H}_{\mathrm{kin},L(R)}(k_{y},k_{z})$ is the kinetic Hamiltonian, $\Delta_{L(R)}(k_{y},k_{z})$ is the superconducting pairing order parameter, and $\phi_{L(R)}$ is the $\mathrm{U}(1)$ superconducting phase, taken to be uniform in the respective superconductor. $c_{k_{y},k_{z};\sigma,n_{x}}$ is the annihilation operator for an electron at site $n_{x}$ with spin $\sigma$ and conserved momentum $k_{y}\hat{y}+k_{z}\hat{z}$.
The tight-binding analogue of the tunneling Hamiltonian for the Josephson junction in Eq. (3) is given by
@@
where $\hat{t}(k_{y},k_{z})=t_{0}(k_{y},k_{z})\sigma_{0}+\mathbf{t}(k_{y},k_{z})\cdot\bm{\sigma}$ describes spin-independent and spin-dependent hopping across the junction barrier. Here, we consider only one site at the barrier for simplicity. Including more sites will increase the amount of in-gap Andreev states but not affect the overall qualitative features of the Josephson current, e.g. its periodicity.
@@ Appendix E Tight-binding models of monopole superconductors
In this section, we outline the tight-binding model of the monopole superconductor described in Eq. (13) used for numerical calculations of monopole superconductors. The general Hamiltonian of a monopole superconductor takes the following form
$H_{\text{MSC}}=$ $\sum_{\mathbf{n},\sigma;\mathbf{n}^{\prime},\sigma^{\prime}}c^{\dagger}_{\mathbf{n},\sigma}[\mathcal{H}_{\mathrm{kin}}]_{\mathbf{n},\sigma;\mathbf{n}^{\prime},\sigma^{\prime}}c_{\mathbf{n}^{\prime},\sigma^{\prime}}$ (84)
$+\sum_{\mathbf{n},\sigma;\mathbf{n}^{\prime},\sigma^{\prime}}\left(c^{\dagger}_{\mathbf{n},\sigma}[\Delta_{\mathrm{inter}}]_{\mathbf{n},\sigma;\mathbf{n}^{\prime},\sigma^{\prime}}c^{\dagger}_{\mathbf{n}^{\prime},\sigma^{\prime}}+\mathrm{h.c.}\right)$
where $\mathcal{H}_{\mathrm{kin}}$ and $\Delta_{\mathrm{inter}}$ are respectively the kinetic and inter-Fermi-surface pairing kernels of the BdG Hamiltonian. Here, $c_{\mathbf{n},\sigma}$ annihilates a single-particle state at lattice site $\mathbf{n}$ and with spin $\sigma=\uparrow,\downarrow$. We take the following convention for the tensor product basis, with the two-dimensional spin basis indexed by $\sigma$ nested under lattice site basis indexed by $\mathbf{n}$.
E.1 Tight-binding model of $q_{p}=-1$ monopole superconductor
For the completeness of the manuscript, we reproduce the tight-binding model of the monopole superconductor with pair monopole charge $q_{p}=-1$ Sun et al. (2019) which corresponds to the continuum model in Eq. (18). The Fourier transform of the tight-binding model of the magnetic doped Weyl semimetal band Hamiltonian on a cubic lattice is given by
$\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}(\mathbf{k})=$ $t_{z}(\cos K_{0,z}-\cos k_{z})\sigma_{z}$ (85)
$+m(\cos k_{x}+\cos k_{y}-2)\sigma_{z}$
$+t_{x}\sin k_{x}\sigma_{x}+t_{y}\sin k_{y}\sigma_{y}-\mu\sigma_{0},$
where $\mu$ is the chemical potential and $t_{i=x,y,z}$ are the hopping amplitudes along the $i^{\text{th}}$ direction (we take $t_{x}=t_{y}=t_{z}\sin K_{0,z}=v_{F}$ to compare to Eq. (19)). At appropriate doping, the system gives rise to two disjoint Fermi surfaces surrounding Weyl nodes at ±$\pm\mathbf{K}_{0}=(0,0,\pm K_{0,z})^{\mathrm{T}}$. The parameter $m$ is chosen satisfying $m/t_{z}\leq-2$ so that the sign of the $\sigma_{z}$-term in momentum space only changes between the two Weyl nodes at ±$\pm\mathbf{K}_{0}$.
@@ E.1 Tight-binding model of $q_{p}=-1$ monopole superconductor
n;n $=(t_{z}\cos K_{0,z}-2m)\sigma_{z}-\mu\sigma_{0},$ (86)
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{x}}$ $=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n}+\bm{\delta}_{x};\mathbf{n}}^{\dagger}=\frac{1}{2}(m\sigma_{z}-it_{x}\sigma_{x}),$
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{y}}$ $=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n}+\bm{\delta}_{y};\mathbf{n}}^{\dagger}=\frac{1}{2}(m\sigma_{z}-it_{y}\sigma_{y}),$
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{z}}$ $=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n}+\bm{\delta}_{z};\mathbf{n}}^{\dagger}=-\frac{1}{2}t_{z}\sigma_{z}.$
Here, $\bm{\delta}_{x}=(1,0,0)^{\mathrm{T}}$, $\bm{\delta}_{y}=(0,1,0)^{\mathrm{T}}$, and $\bm{\delta}_{z}=(0,0,1)^{\mathrm{T}}$ are lattice unit vectors along the crystalline axes. The $s$-wave pairing in Eq. (84) is defined locally on-site as
@@
where $\Delta_{0}$ is the pairing amplitude.
@@
Under open boundary conditions, the system exhibits helical surface states. Consider the partial Fourier-transform of the above tight-binding model, for given conserved $\tilde{k}_{z}$, the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole pairing order can be regarded as an effective two-dimensional chiral pairing order characterized by local angular momentum $l_{z,\mathrm{loc}}=\pm 1$. Away from $\mathrm{FS}_{1}$, the chiral surface modes result from the Weyl Fermi arcs; however, between the gap nodes at the poles of $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$, a single surface mode appears within the gap. The single surface mode appearing between the gap nodes is a result of the effective local chiral $p$-wave pairing, as described in Sec. V. As $\tilde{k}_{z}$ varies from the south to north pole, corresponding to Fig. 9 (b-d), though the topology of the pairing order, which is given by pair monopole charge $q_{p}=-1$, does not change, the sign of the in-gap states’ group velocity changes. The change in the group velocity is due to the shift in the local angular momentum, in agreement to the effective pairing order in Eq. (24).
@@
Figure 9: (a) Fermi surface $\mathrm{FS}_{1}$ of the tight-binding model $\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}$ in Appendix E.1 (black, solid) at $\tilde{k}_{x}=0$. Three dashed lines correspond to three values of relative momenta, $\tilde{k}_{z}=-0.20$, $0.00$, and $0.17$. (b)-(d) The bulk (gray) and surface (colored) BdG spectra of the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor with boundary at $x=0$ and at the three values of $\tilde{k}_{z}$ shown in (a). The color scheme indicates the charge of the surface BdG quasiparticles states, with particle-like (hole-like) states corresponding to red (blue). We use parameters $t_{x}=t_{y}=t_{z}\sin K_{0}=-0.5$, $K_{0}=1.0$, $\mu=-0.1$, $m=1.0$, and $\Delta_{0}=0.02$.
@@
The bulk monopole pairing order of the tight-binding model is shown in Fig. 10. The pairing between two helical Fermi surfaces $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$, as introduced in Sec. III.1, is defined as $\Delta_{\mathrm{MSC}}^{(q_{p})}(\tilde{\mathbf{k}})=\langle\chi_{1,+}(\tilde{\mathbf{k}})|\Delta_{\mathrm{inter}}(\mathbf{k})|\chi_{2,+}^{*}(-\tilde{\mathbf{k}})\rangle$ in Eq. (16). Here, $|\chi_{1(2),+}(\tilde{\mathbf{k}})\rangle$ is the helical eigenstate of $\mathrm{FS}_{1(2)}$. The magnitude of the effective pairing is shown in Fig. 10(a) and 10(b), where there are two gap nodes pinned to the north and south pole of $\mathrm{FS}_{1}$. The phase winding, chosen in a gauge where the pairing is analytic at the north and south pole, is shown in Fig. 10(c) and (d) respectively. Near the south pole, consistent with the group velocity of the surface modes in Fig. 9(b), the monopole pairing order behaves locally as a $p_{x}+ip_{y}$ chiral pairing order. In contrast, near the north pole, the effective pairing order behaves locally as a $p_{x}-ip_{y}$ chiral pairing order. The shift in local angular momentum $l_{z,\mathrm{loc}}=\pm 1$ from the global angular momentum, $l_{z,\mathrm{glob}}=0$, is indicative of the nontrivial pair monopole charge, $q_{p}=-1$. For the tight-binding model, the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconducting pairing order can be expressed as a superposition of monopole harmonics, all which have pair monopole charge $q_{p}=-1$ and global angular momentum $l_{z}=2$, but reside in different partial wave channels ($l\geq|q_{p}|$) due to the lattice symmetry.